OpenStudy (anonymous):

I am stuck with a statistics problem based on exponential distribution, can anyone help?

OpenStudy (anonymous):

Hey Arjun - maybe... What's your problem?

OpenStudy (anonymous):

Post the problem i might be able to help out

OpenStudy (anonymous):

Thanks for the response folks. The problem is: After given an antibiotic in dogs, certain type of bacteria dies completely within six to twelve hours based on an exponential distribution; i.e. f(x) = 0.95 e ^{-x/3}., where \[6\le x \le\] . What is the probability that a dog receiving the antibiotic will be completely free of bacteria in less than eight hours?

OpenStudy (anonymous):

x is greater than and equal to 6, and x is less than and equal to 12

OpenStudy (waheguru):

wow hard

OpenStudy (anonymous):

you could integrate the function and evaluate first on the interval 6 to 12 and then from 6 to 8. and do the following:\[\int\limits_{6}^{8}f(x) \div \int\limits_{6}^{12}f(x)\] it's just a thought, i'm not entirely sure

OpenStudy (anonymous):

Thanks estebananaya, m not too completely sure but trying right now

OpenStudy (turingtest):

I knew Zarkon would shoot over here. I'm posting so I can see the solution if you get it.

OpenStudy (anonymous):

Well that would be great if Zarkon can figure this out..

OpenStudy (zarkon):

the pdf you gave doesn't integrate to 1.

OpenStudy (anonymous):

so, do you think there is a typo in question.. I tried estebananaya's way (posted above) and it came down to .25, but not sure if its right..

OpenStudy (zarkon):

if \(f\) is a true pdf then it is required that \[\int_{\Omega}f(x)dx=1\] where \(\Omega\) is the support of the distribution

OpenStudy (anonymous):

well, no doubts about that.. and one other way I see it, is to subtract individual probabilities, I mean calculating at 8 and 6, and subtracting.. but not sure..

OpenStudy (zarkon):

I would double check the problem to make sure you have typed it correctly. If you have typed it correctly then the problem is ill posed since they didn't give you a true density function

OpenStudy (anonymous):

Zarkon, in question there is an incomplete equation for x, it has to be \[6\le x \le12\]

OpenStudy (zarkon):

for your \(f\) \[\int\limits_{6}^{12}f(x)dx\ne 1\]

OpenStudy (anonymous):

thanks